OpenStudy (anonymous):

some one please help evaluate -5g-6for g=-2 a.4 b.16 c.10 d.-13

Directrix (directrix):

-5g - 6 for g=-2 -------------------------- = - 5 * (-2) - 6 =

Directrix (directrix):

@kate_bell Any idea

OpenStudy (anonymous):

I got it thanks

OpenStudy (anonymous):

Replace g with negative two and then youll have -5(-2)-6 then just use a calculator :)

Directrix (directrix):

What did you get for an answer? We can check it.

OpenStudy (anonymous):

4

Directrix (directrix):

- 5 * (-2) - 6 = 10 - 6 = 4 That's what I got, too.

OpenStudy (anonymous):

lol k:)

OpenStudy (anonymous):

thanks